# Quantum optical coherence theory based on Feynman’s path integral

### v1

@@ IV Second-order interference of light
As mentioned above, nearly all the studies of optical coherence are about the first-order interference of light before 1956 Born and Wolf (2013). In 1956, Hanbury Brow and Twiss discovered two-photon bunching of thermal light Brown and Twiss (1956b, a), which is one of the most famous second-order interference phenomena in quantum optics. The history of second-order coherence of light can trace back to 1940s when several groups measured the correlation of photon pair generated in the annihilation of a positron-electron pair Hanna (1948); Bleuler and Bradt (1948); Wu and Shaknov (1950). The second-order interference of light is measured with coincidence count detection system, which is proposed by Bothe in 1925 Bothe and Geiger (1925).
@@
The second-order interference of light has an unique role in the development of optical coherence theory. The first-order interference of light exists in nature. For instance, soap bubbles, transparent wings of insects, some butterfly wings, etc. appear colored streaks when exposed to sunlight is due to the first-order interference of light. The second-order interference of light does not exist in nature. It is a human invented phenomenon. Special detection scheme is required to observe the second-order interference of light. The second-order coherence is qualitatively different from the first-order interference. It may be reason why people took more than 200 years since the first observed phenomenon of first-order interference of light to the first observed phenomenon of second-order interference of light. Extending the second-oder coherence to the third- and higher-order coherence is natural. The third- and higher-order coherence is quantitative different from the second-order coherence.
@@
Figure 10: (a) First-, (b) second- and (c) third-order interference. S is light source. IT1, IT2, IT3 are first-, second- and third-order interferometers, respectively. D is photon detector. CC is coincidence count detection scheme.
@@
Figure 10 shows the schemes for the first-, second- and third-order interference of light. For comparison, the first-order interference of light shown in Fig. 10(a) is illustrated first. In a typical experiment of first-order interference of light, light generated by the light source S is split into more than one components in the first-order interferometer IT1. Then the split light beams are recombined to observe the intensity distribution over space or time by a single photo-detector. Young’s double-slit interferometer, Mach-Zehnder interferometer, Michelson interferometer are the typical first-order interferometers Born and Wolf (2013). The first-order interference of two and multiple independent light beams introduced in Sects. III.2 and III.3 belong to the first-order interference of light.
@@
A typical scheme for the second-order interference of light is shown in Fig. 10(b). Light emitted by source S is split into two or more different components in the second-order interferometer (IT2). Then the output light beams of IT2 are detected by two detectors (D1 and D2). Hanbury Brown-Twiss (HBT) interferometerBrown and Twiss (1956b, a), Hong-Ou-Mandel (HOM) interferometer Hong et al. (1987); Shih and Alley (1988); Alley and Shih (1986), Franson interferometer Franson (1989) are some typical second-order interferometers. The output signals of these two detectors are sent into coincidence count detection system (CC) to measure the correlation of these two signals. Based on what types of photo-detector is employed, there are two work principles for CC. The first one is for normal photo-electric detector. The output of the detector is electric current or voltage proportional to the intensity of the incidental light. The measured correlation is
@@
where $I_{1}(\vec{r}_{1},t_{1})$ is the intensity of incidental light at space-time coordinate $(\vec{r}_{1},t_{1})$ detected by D1, $I_{2}(\vec{r}_{2},t_{2})$ is the intensity of incidental light at space-time coordinate $(\vec{r}_{2},t_{2})$ detected by D2, delimited-⟨$\langle$..$\rangle$ … ⟩ is ensemble average. For ergodic and stationary system, ensemble average can be approximated by time average Mandel and Wolf (1995). Equation (47) measures the correlation of intensity fluctuations detected by these two detectors. The CC system based on normal photo-electric detector was employed by Hanbury Brown and Twiss in their correlation experiments Brown and Twiss (1956b, a).
The second one is for single-photon detector. The output of single-photon detector is a voltage pulse. Equation (47) can not be employed to calculate the correlation of the outputs from single-photon detectors. A different coincidence count principle should be employed. When the time difference between these two photon detection events is less than a time window $T$, these two photon detection events is a two-photon coincidence count. Otherwise, these two photon detection is not a two-photon coincidence count. The time window $T$ should be less than the coherence time due to these two photons are indistinguishable within the same coherence volume. There are also other coincidence count detection system, such as two-photon absorption Boitier et al. (2009), sum frequency generation Pe’er et al. (2007), which are also based on the same principle as the one for single-photon detector.
The second-order interference shown in Fig. 10(b) can be extended to the third- and higher-order interference. Figure 10(c) shows the scheme for the third-order interference of light. The light beam emitted by source S is split into multiple components and detected by three detectors, D1, D2 and D3. The output of these three detectors are sent into a three-photon coincidence count system (CC) to measure the third-order correlation. Based on different types of detectors are employed, there are also two different types of principle for three-photon coincidence count system. When normal photo-detectors are employed, the third-order correlation is given by
@@
where the meanings of the symbols in Eq. (48) are similar as the ones in Eq. (47). When single-photon detectors are employed, the time differences between these three photon detection events are less than a time window $T$, the three photon detection events is a three-photon coincidence count. Otherwise, the three photon detection events is not a three-photon coincidence count. Similar method can be employed to define the fourth- and higher-order interference of light.
@@ IV.1 Second-order interference of one light beam
Figure 11 shows the second-order interference one light beam in a HBT interferometer. The light beam emitted by source S is split into two by a 1:1 non-polarizing beam splitter BS. The output light beams of BS are detected by two single-photon detectors, D1 and D2. The outputs of these two detectors are sent into two-photon coincidence count system CC to measure the second-order correlation of light.
Figure 11: One light beam in a HBT interferometer. S is light source. BS is 1:1 nonpolarizing beam splitter. D1 and D2 are two single-photon detectors. CC is two-photon coincidence count system.
@@ IV.1.1 Second-order interference of one thermal light beam
Assuming S is a thermal light source in Fig. 11, the scheme shown in Fig. 11 is equivalent to the scheme employed by Hanbury Brown and Twiss discovering two-photon bunching of thermal light in 1956 Brown and Twiss (1956b, a). There are two different paths for D1 and D2 detecting a two-photon coincidence count in Fig. 11. The first path is photon a is detected by D1 and photon b is detected by D2. The probability for the path is $1/2$ and the corresponding two-photon probability amplitude is $A(a,D_{1};b,D_{2})$. The second path is photon a is detected by D2 and photon b is detected by D1. The probability for the path is $1/2$ and the corresponding two-photon probability amplitude is $A(a,D_{2};b,D_{1})$. When photons a and b are in the same coherence volume of thermal light, these two different paths to trigger a two-photon coincidence count are indistinguishable. The two-photon probability function for the $j$th detected photon pair is Feynman et al. (2010, 2011)
@@
$=$ $|A_{j}(a,D_{1};b,D_{2})+A_{j}(a,D_{2};b,D_{1})|^{2}$
$=$ $|e^{i(\varphi_{aj}+\varphi_{bj})}K_{j}(a,D_{1};b,D_{2})$
@@
where $\varphi_{aj}$ and $\varphi_{bj}$ are the initial phases of photons a and b in the $j$th detected photon pair, respectively, $K_{j}(a,D_{1};b,D_{2})$ is photon’s Feynman propagator for photon a emitted by S going to D1 and photon b emitted by S going to D2. The meaning of $K_{j}(a,D_{2};b,D_{1})$ is defined similarly. Based on the principle for the probability amplitude of independent events introduced by Feynman, $K_{j}(a,D_{1};b,D_{2})=K_{j}(a,D_{1})K_{j}(b,D_{2})$. Equation (49) can be simplified as
@@
The two-photon probability function of $N$ detected photon pairs is the sum of the probabilities of $N$ detected photon pairs,
@@
is employed to eliminate the influence of the first-order interference on the second-order interference, where $P^{(1)}(\vec{r}_{1},t_{1})$ and $P^{(1)}(\vec{r}_{2},t_{2})$ are single-photon probability function at space-time coordinates $(\vec{r}_{1},t_{1})$ and $(\vec{r}_{2},t_{2})$, respectively. $P^{(1)}(\vec{r}_{1},t_{1})$ is written as $P(\vec{r}_{1},t_{1})$ if no misunderstanding exists. Two-photon probability function $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ and normalized two-photon probability function $p^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ correspond to the second-order coherence function $G^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ and normalized second-order coherence function $g^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ in Glauber’s quantum optical coherence theory, respectively Glauber (1963a, c).
Let us first calculate temporal two-photon bunching of thermal light. Assuming S is a point light source, D1 and D2 are at symmetrical positions of BS to simplify the calculation. In this case, photon’s Feynman propagator is
@@
where the spatial part is ignored, $\omega_{0}$ is the central angular frequency, $\Delta\omega$ is angular frequency distribution function of thermal light, $t_{1}$ is the time of photon detection event at D1, $t_{a}$ is the emitting time of photon a. Substituting Eq. (53) into Eq. (IV.1.1) and considering $f(\omega)$ equals 1 within $[\omega_{0}-\Delta\omega/2,\omega_{0}+\Delta\omega/2]$ and equals 0 in other domains, the temporal two-photon probability function is Shih (2020)
@@
It is easy to prove that $P^{(1)}(t_{1})$ and $P^{(2)}(t_{1})$ are constant for thermal light. The normalized temporal two-photon probability function, $p^{(2)}(t_{1}-t_{2})$ is proportional to $P^{(2)}(t_{1}-t_{2})$. In the following, if the normalized probability function is proportional to the corresponding probability function, only probability function is calculated.
Figure 12: Temporal two-photon bunching of thermal light. $t_{1}-t_{2}$ is the relative position difference between D1 and D2. Lines 1, 2, 3 correspond to the frequency bandwidths equal $\Delta\omega$, $2\Delta\omega$, and $4\Delta\omega$, respectively.
Figure 12 shows the temporal two-photon probability functions of thermal light with different frequency bandwidths. When two detectors are at symmetrical positions, the probability of detecting two photons equal 2. As the time difference, $|t_{1}-t_{2}|$, increases, the probability of detecting two photons decreases. When the time difference between two photon detection events is much longer than the coherence time, the probability of detecting two photons equal 1, which means that these two photon detection events are independent.
The spatial two-photon bunching of thermal light can be calculated in the similar way. Assuming one dimension light source is employed, the light emitted by S is single frequency light, and the distance between the source plane and D1 is equal the one of source plane and D2. photon’s Feynman propagator in this case can be approximated by Green function in classical optics Goodman (2005)
@@
where $d$ is the length of thermal light source, $s(x_{s})$ is the distribution function of S, $\vec{r}_{1}$ is position vector of D1, $\vec{r}_{s}$ is the position vector of S emitting photon a, $\vec{k}_{s1}$ is wave vector of light. Substituting Eq. (55) into Eq. (IV.1.1) and considering $s(x_{s})$ equals 1 within $[-d/2,d/2]$ and equals 0 in other domains, spatial two-photon probability function of thermal light in a HBT interferometer is Shih (2020)
@@
Based on the results above, two-photon bunching of thermal light can be interpreted as the result of two-photon interference, which is the superposition of two-photon probability amplitudes for two different and indistinguishable paths to trigger a two-photon coincidence count Fano (1961); Feynman (2006); Glauber (1995). Two-photon interference in the second-order interference of light can not be used to disprove Dirac’s statement that different photons never interferes with each other in the first-order interference of light Paul (1986). By analogy, Dirac’s statement about the first-order interference of light can be extended to the second-order interference of light: Each photon pair only interferes with itself and interferes between two different photon pairs never occurs Shih (2020).
@@ IV.1.2 Second-order interference of one superbunching light beam
$g^{(2)}(0)$ equals 2 for thermal light. $g^{(2)}(0)$ measured in the HBT interferometer is the degree of second-order coherence of light, which is usually employed to categorize different types of light Loudon (2000). $g^{(2)}(0)$ equals 1 for single-mode continuous-wave laser light. When $g^{(2)}(0)$ is less than 1, the light belongs to non-classical light. When $g^{(2)}(0)$ is larger than 2, it is superbunching light Ficek and Swain (2005). Superbunching light can be either classical or nonclassical. For instance, entangled photon pairs is non-classical superbunching light Klyshko (1996) and superbunching pseudothermal light is classical light Zhou et al. (2017). This section will concentrate on the second-order interference of a classical superbunching light beam in HBT interferometer.
There are two different methods to increase the degree of second-order coherence of classical light. The first one is adding more two-photon paths. The second one is the two-photon probability amplitudes changes with time.
Figure 13: Increase the degree of second-order coherence by adding more two-photon paths.
Figure 13 shows how to increase the degree of second-order coherence by adding more two-photon paths. Figure 13(a) shows there are two different ways for photon a and b to trigger a two-photon coincidence count in D1 and D2. The first one is photon a is detected by D1 and photon b is detected by D2. The corresponding two-photon probability amplitude is $A(a,D_{1};b,D_{2})$. The second one is photon a is detected by D2 and photon b is detected by D1. The corresponding two-photon probability amplitude is $A(a,D_{2};b,D_{1})$. If these two different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
@@
$=$ $\langle|A(a,D_{1};b,D_{2})+A(a,D_{2};b,D_{1})|^{2}\rangle.$
@@
$=$ $\langle|A(a,D_{1};b,D_{2})|^{2}\rangle+\langle|A(a,D_{2};b,D_{1})|^{2}\rangle,$
which will contribute to the background of two-photon probability function. Figure 13(a) and Eqs. (57) are valid for two-photon bunching of thermal light. $g^{(2)}(0)$ equals 2 for thermal light. There is another way to estimate the value of $g^{(2)}(0)$. Assuming all the photons have equal probability reaching D1 and D2, $|A(a,D_{1};b,D_{2})|$ equals $|A(a,D_{2};b,D_{1})|$. The are four terms in Eq. (57) and $P^{(2)}(0)$ equals $4|A|^{2}$, where $A$ is short for $|A(a,D_{1};b,D_{2})|$. There are 2 terms in Eq. (58) and $P^{(2)}(0)$ equals $2|A|^{2}$. The degree of second-order coherence equals the normalized two-photon probability function when these two detectors are at the same space-time coordinate,
$g^{(2)}(0)=p^{(2)}(0)=\frac{4|A|^{2}}{2|A|^{2}}=2.$ (59)
The reason why Eq. (58) can be treated as the background is that these two paths are distinguishable. The two photon detection events are independent and $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ equals $P^{(1)}(\vec{r}_{1},t_{1})P^{(1)}(\vec{r}_{2},t_{2})$.
Similar method can be employed to calculate the degree of second-order coherence of light in Fig. 13(b). There are four different paths for photons a and b triggering a two-photon coincidence count. The first one is photon a transmitting through $\alpha$ is detected by D1 and photon b transmitting through $\beta$ is detected by D2. The corresponding two-photon probability amplitude is $A(a,\alpha,D_{1};b,\beta,D_{2})$. The other three paths can be defined similarly and the corresponding two-photon probability amplitudes are $A(a,\alpha,D_{2};b,\beta,D_{1})$, $A(a,\beta,D_{1};b,\alpha,D_{2})$ and $A(a,\beta,D_{2};b,\alpha,D_{1})$. When these four different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function
@@
$=$ $\langle|A(a,\alpha,D_{1};b,\beta,D_{2})+A(a,\alpha,D_{2};b,\beta,D_{1})$
$+A(a,\beta,D_{1};b,\alpha,D_{2})+A(a,\beta,D_{2};b,\alpha,D_{1})|^{2}\rangle.$
When these four different paths are distinguishable, the two-photon probability function
@@
$=$ $\langle|A(a,\alpha,D_{1};b,\beta,D_{2})|^{2}\rangle+\langle|A(a,\alpha,D_{2};b,\beta,D_{1})|^{2}\rangle$
$+\langle|A(a,\beta,D_{1};b,\alpha,D_{2})|^{2}\rangle+\langle|A(a,\beta,D_{2};b,\alpha,D_{1})|^{2}\rangle$
Again, assuming photons have equal probability being detected by D1 and D2. The ratio between Eqs. (IV.1.2) and (IV.1.2) at the same space-time coordinate equals the degree of second-order coherence,
$g^{(2)}(0)=p^{(2)}(0)=\frac{16|A^{\prime}{}|^{2}}{4|A^{\prime}{}|^{2}}=4,$ (62)
@@
Figure 14: Experimental scheme to observe two-photon superbunching with $2N$ different paths Zhou et al. (2017). P1, P2, …, PN are $N$ pinholes. RG1, RG2, …, RGN are $N$ rotating groundglasses.

### v2

@@ IV Second-order interference of light
As mentioned above, nearly all the studies of optical coherence are about the first-order interference of light before 1956 Born and Wolf (2013). In 1956, Hanbury Brown and Twiss discovered two-photon bunching of thermal light Brown and Twiss (1956b, a), which is regarded as one of the cornerstones of modern quantum optics. The history of second-order coherence of light can trace back to 1940s when several groups measured the correlation of photon pair generated in the annihilation of positron-electron pairs Hanna (1948); Bleuler and Bradt (1948); Wu and Shaknov (1950). The key element to measure the second-order interference of light is coincidence count detection system, which is proposed by Bothe in 1925 Bothe and Geiger (1925).
@@
The second-order interference of light plays an unique role in the history of the development of optical coherence. The first-order interference of light exists in nature. For instance, soap bubbles, transparent wings of insects, some butterfly wings, etc. appearing colored streaks when exposed to sunlight is due to the first-order interference of light. The second-order interference of light does not exist in nature. It is a human-invented phenomenon. Special detection scheme is required to observe the second-order interference of light. The second-order interference of light is qualitatively different from the first-order interference of light. It may be the reason why people took more than 200 years to observe the first phenomenon of the second-order interference of light since the first observation of the first-order interference of light. On the other hand, generalizing the second-oder interference of light to the third- and higher-order interference of light is straightforward. The third- and higher-order interference of light is quantitatively different from the second-order interference of light.
@@
Figure 10: (a) First-, (b) second- and (c) third-order interference of light. S is a light source. IT1, IT2, and IT3 are first-, second- and third-order interferometers, respectively. D is a single-photon detector. CC is a coincidence count detection scheme.
@@
Figure 10 shows the schemes for the first-, second- and third-order interference of light. For comparison, the first-order interference of light shown in Fig. 10(a) is illustrated first. In a typical experiment of first-order interference of light, light generated by the light source, S, is split into more than one components in the first-order interferometer, IT1. Then the split light beams are recombined to observe the intensity distribution over space or time by one detector. Young’s double-slit interferometer, Mach-Zehnder interferometer, Michelson interferometer are some typical first-order interferometers Born and Wolf (2013). The first-order interference of two and multiple independent light beams introduced in Sects. III.2 and III.3 belong to the first-order interference of light.
@@
A typical scheme for the second-order interference of light is shown in Fig. 10(b). Light emitted by S is split into two or more different components in the second-order interferometer, IT2. Then the output light beams of IT2 are detected by two detectors, D1 and D2. Hanbury Brown-Twiss (HBT) interferometer Brown and Twiss (1956b, a), Hong-Ou-Mandel (HOM) interferometer Hong et al. (1987); Shih and Alley (1988); Alley and Shih (1986), Franson interferometer Franson (1989) are some typical second-order interferometers. The output signals of these two detectors are sent into a coincidence count detection system, CC, to measure the correlation of these two signals. Based on what types of detector is employed, there are two work principles for CC. The first one is for normal detectors. The output of the detector is electric current or voltage proportional to the intensity of the incidental light. The measured correlation is given by the product of these two output signals,
@@
where $I_{1}(\vec{r}_{1},t_{1})$ is the intensity of incidental light at space-time coordinate $(\vec{r}_{1},t_{1})$ detected by D1, $I_{2}(\vec{r}_{2},t_{2})$ is the intensity of incidental light at space-time coordinate $(\vec{r}_{2},t_{2})$ detected by D2, delimited-⟨$\langle$..$\rangle$ … ⟩ is ensemble average. For ergodic and stationary system, ensemble average can be approximated by time average Mandel and Wolf (1995). Equation (47) measures the correlation of intensity fluctuations detected by these two detectors. The CC system based on normal photo-electric detectors was employed by Hanbury Brown and Twiss in their experiments Brown and Twiss (1956b, a).
The second principle is for single-photon detectors. The output of a single-photon detector is a voltage pulse. Equation (47) can not be employed to calculate the correlation of the outputs from two single-photon detectors. A different coincidence count principle should be employed. When the time difference between these two photon detection events is less than a time window $T$, these two photon detection events are a two-photon coincidence count. Otherwise, these two photon detection events are not a two-photon coincidence count. The time window $T$ should be less than the coherence time due to these two photons are indistinguishable within the same coherence volume. There are also other coincidence count detection system, such as two-photon absorption Boitier et al. (2009), sum frequency generation Pe’er et al. (2007), which are also based on the same principle as the one for single-photon detectors.
The second-order interference of light shown in Fig. 10(b) can be extended to the third- and higher-order interference of light. Figure 10(c) shows the scheme for the third-order interference of light. The light beam emitted by S is split into multiple components and detected by three detectors, D1, D2 and D3. The output of these three detectors are sent into a three-photon coincidence count system, CC, to measure the third-order correlation. Based on different types of the employed detectors, there are also two different principles for three-photon coincidence count system. When normal detectors are employed, the third-order correlation is given by
@@
where the meanings of the symbols in Eq. (48) are similar as the ones in Eq. (47). When single-photon detectors are employed, the time differences between these three photon detection events are less than a time window $T$, the three photon detection events are a three-photon coincidence count. Otherwise, the three photon detection events are not a three-photon coincidence count. Similar method can be employed to define the fourth- and higher-order interference of light.
@@ IV.1 Second-order interference of one light beam
Figure 11 shows the scheme for the second-order interference one light beam in a HBT interferometer. The light beam emitted by source S is split into two by a 1:1 non-polarizing beam splitter, BS. The output light beams of BS are detected by two single-photon detectors, D1 and D2. The outputs of these two detectors are sent into a two-photon coincidence count detection system, CC, to measure the second-order correlation of light.
Figure 11: Second-order interference of one light beam in a HBT interferometer. S is a light source. BS is a 1:1 non-polarizing beam splitter. D1 and D2 are two single-photon detectors. CC is a two-photon coincidence count detection system.
@@ IV.1.1 Second-order interference of one thermal light beam
Assuming S is a thermal light source in Fig. 11, the scheme shown in Fig. 11 is equivalent to the scheme employed by Hanbury Brown and Twiss in 1956 Brown and Twiss (1956b, a). There are two different paths for D1 and D2 detecting a two-photon coincidence count. The first path is that photon a is detected by D1 and photon b is detected by D2. The probability for the path is $1/2$ and the corresponding two-photon probability amplitude is $A(a,D_{1};b,D_{2})$. The second path is that photon a is detected by D2 and photon b is detected by D1. The probability for the path is $1/2$ and the corresponding two-photon probability amplitude is $A(a,D_{2};b,D_{1})$. When photons a and b are in the same coherence volume of thermal light, these two different paths to trigger a two-photon coincidence count are indistinguishable. The two-photon probability function for the $j$th detected photon pair is Feynman et al. (2010, 2011)
@@
$=$ $|\frac{1}{\sqrt{2}}[A_{j}(a,D_{1};b,D_{2})+A_{j}(a,D_{2};b,D_{1})]|^{2}$
$\propto$ $|e^{i(\varphi_{aj}+\varphi_{bj})}K_{j}(a,D_{1};b,D_{2})$
@@
where $\varphi_{aj}$ and $\varphi_{bj}$ are the initial phases of photons a and b in the $j$th detected photon pair, respectively, $K_{j}(a,D_{1};b,D_{2})$ is photon’s Feynman propagator for photon a emitted by S going to D1 and photon b emitted by S going to D2. The meaning of $K_{j}(a,D_{2};b,D_{1})$ is defined similarly. Based on the principle for calculating the probability amplitude of independent events introduced by Feynman Feynman et al. (2011), $K_{j}(a,D_{1};b,D_{2})$ equals $K_{j}(a,D_{1})K_{j}(b,D_{2})$. Equation (49) can be simplified as
@@
The two-photon probability function of $N$ detected photon pairs is the sum of the probability functions of $N$ detected photon pairs,
@@
is employed to eliminate the influence of the first-order interference pattern on the second-order interference pattern, where $P^{(1)}(\vec{r}_{1},t_{1})$ and $P^{(1)}(\vec{r}_{2},t_{2})$ are single-photon probability functions at space-time coordinates $(\vec{r}_{1},t_{1})$ and $(\vec{r}_{2},t_{2})$, respectively. $P^{(1)}(\vec{r}_{1},t_{1})$ is usually written as $P(\vec{r}_{1},t_{1})$ for short if no possible misunderstanding exists. The two-photon probability function, $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$, and the normalized two-photon probability function, $p^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$, correspond to the second-order coherence function, $G^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$, and the normalized second-order coherence function, $g^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$, in Glauber’s quantum optical coherence theory, respectively Glauber (1963a, c).
Let us first calculate temporal two-photon bunching of thermal light. Assume S is a point thermal light source. D1 and D2 are at symmetrical positions of BS to simplify the calculation. In this case, photon’s Feynman propagator is
@@
where the spatial part is ignored, $\omega_{0}$ is the central angular frequency, $\Delta\omega$ is the angular frequency distribution function of thermal light, $t_{1}$ is the time of photon detection event at D1, $t_{a}$ is the emitting time of photon a. Substituting Eq. (53) into Eq. (IV.1.1) and assuming $f(\omega)$ equals 1 within $[\omega_{0}-\Delta\omega/2,\omega_{0}+\Delta\omega/2]$ and equals 0 in other domains, the temporal two-photon probability function is Shih (2020)
@@
It is easy to prove that $P^{(1)}(t_{1})$ and $P^{(1)}(t_{2})$ are constant for thermal light. The normalized temporal two-photon probability function, $p^{(2)}(t_{1}-t_{2})$ is proportional to $P^{(2)}(t_{1}-t_{2})$. In the following parts, if the normalized probability function is proportional to the corresponding probability function, only probability function is calculated.
Figure 12: Temporal two-photon bunching of thermal light. $t_{1}-t_{2}$ is the time difference between photon detection events at D1 and D2. Lines 1, 2, 3 correspond to the frequency bandwidths equaling $\Delta\omega$, $2\Delta\omega$, and $4\Delta\omega$, respectively.
Figure 12 shows the temporal two-photon probability functions of thermal light with different frequency bandwidths. When $t_{1}$ equals $t_{2}$, the probability of detecting two photons equals 2. As the time difference, $|t_{1}-t_{2}|$, increases, the probability of detecting two photons decreases. When the time difference between two photon detection events is much longer than the coherence time, the probability of detecting two photons equals 1, which indicates that these two photon detection events are independent.
The spatial two-photon bunching of thermal light can be calculated in the same way. Assume one dimension thermal light source is employed. The light emitted by S is single-frequency light. The distance between the source plane and D1 equals the one between source plane and D2. The photon’s Feynman propagator in this case can be approximated by Green function in classical optics Goodman (2005)
@@
where $d$ is the length of thermal light source, $s(x_{s})$ is the intensity distribution function of S, $\vec{r}_{1}$ is the position vector of D1, $\vec{r}_{s}$ is the position vector of S emitting photon a, $\vec{k}_{s1}$ is the wave vector of light. Substituting Eq. (55) into Eq. (IV.1.1) and assuming $s(x_{s})$ equals 1 within $[-d/2,d/2]$ and equals 0 in other domains, spatial two-photon probability function of thermal light in a HBT interferometer is Shih (2020)
@@
Based on the results above, two-photon bunching of thermal light can be interpreted as the result of two-photon interference. It is the result of superposing two-photon probability amplitudes for two different and indistinguishable paths to trigger a two-photon coincidence count Fano (1961); Feynman (2006); Glauber (1995). Two-photon interference in the second-order interference of light can not be used to disprove Dirac’s statement that different photons never interferes with each other, which was meant for the first-order interference of light Paul (1986). By analogy, Dirac’s statement about the first-order interference of light can be generalized to the second-order interference of light: Each photon pair only interferes with itself and interferes between two different photon pairs never occurs Shih (2020).
@@ IV.1.2 Second-order interference of one superbunching light beam
$g^{(2)}(0)$ measured in a HBT interferometer is the degree of second-order coherence of light, which is usually employed to categorize different types of light Loudon (2000). $g^{(2)}(0)$ equals 1 for single-mode continuous-wave laser light. When $g^{(2)}(0)$ is less than 1, the light belongs to non-classical light. $g^{(2)}(0)$ equals 2 for thermal light. When $g^{(2)}(0)$ is larger than 2, it belongs to superbunching light Ficek and Swain (2005). Superbunching light can be either classical or non-classical. For instance, entangled photon pairs is non-classical superbunching light Klyshko (1996) and superbunching pseudothermal light is classical light Zhou et al. (2017). This section will concentrate on the second-order interference of a classical superbunching light beam in a HBT interferometer.
There are two different methods to increase the degree of second-order coherence of classical light. The first one is adding more two-photon paths. The second one is the two-photon probability amplitudes changing with time.
Figure 13: Increase the degree of second-order coherence by adding more two-photon paths. a and b are two photons. D1 and D2 are two single-photon detectors. $\alpha$ and $\beta$ are two nodes for photons.
Figure 13 shows how to increase the degree of second-order coherence by adding more two-photon paths. Figure 13(a) shows there are two different paths for photons a and b to trigger a two-photon coincidence count in D1 and D2. The first path is that photon a is detected by D1 and photon b is detected by D2. The corresponding two-photon probability amplitude is $A(a,D_{1};b,D_{2})$. The second path is that photon a is detected by D2 and photon b is detected by D1. The corresponding two-photon probability amplitude is $A(a,D_{2};b,D_{1})$. If these two different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
@@
$=$ $\langle|\frac{1}{\sqrt{2}}[A(a,D_{1};b,D_{2})+A(a,D_{2};b,D_{1})]|^{2}\rangle.$
@@
$=$ $\frac{1}{2}[\langle|A(a,D_{1};b,D_{2})|^{2}\rangle+\langle|A(a,D_{2};b,D_{1})|^{2}\rangle],$
which consists of the background of two-photon probability function. Figure 13(a) and Eqs. (57) are valid for thermal light in a HBT interferometer. $g^{(2)}(0)$ equals 2 for thermal light. There is another way to estimate the value of $g^{(2)}(0)$. Assuming all the photons have equal probability reaching D1 and D2, $|A(a,D_{1};b,D_{2})|$ equals $|A(a,D_{2};b,D_{1})|$. The are four terms in Eq. (57) and $P^{(2)}(0)$ equals $2|A|^{2}$, where $A$ is short for $|A(a,D_{1};b,D_{2})|$. There are 2 terms in Eq. (58) and $P^{(2)}(0)$ equals $|A|^{2}$. The degree of second-order coherence equals the normalized two-photon probability function when these two detectors are at the same space-time coordinate,
$g^{(2)}(0)=p^{(2)}(0)=\frac{2|A|^{2}}{|A|^{2}}=2.$ (59)
The reason why Eq. (58) can be treated as the background is that these two paths are distinguishable. The two photon detection events are independent and $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ equals $P^{(1)}(\vec{r}_{1},t_{1})P^{(1)}(\vec{r}_{2},t_{2})$ in this case.
Similar method can be employed to calculate the degree of second-order coherence of light in Fig. 13(b). There are four different paths for photons a and b to trigger a two-photon coincidence count. The first path is that photon a transmitting through $\alpha$ is detected by D1 and photon b transmitting through $\beta$ is detected by D2. The corresponding two-photon probability amplitude is $A(a,\alpha,D_{1};b,\beta,D_{2})$. The other three paths can be defined similarly and the corresponding two-photon probability amplitudes are $A(a,\alpha,D_{2};b,\beta,D_{1})$, $A(a,\beta,D_{1};b,\alpha,D_{2})$ and $A(a,\beta,D_{2};b,\alpha,D_{1})$. When these four different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
@@
$=$ $\langle|\frac{1}{\sqrt{4}}[A(a,\alpha,D_{1};b,\beta,D_{2})+A(a,\alpha,D_{2};b,\beta,D_{1})$
$+A(a,\beta,D_{1};b,\alpha,D_{2})+A(a,\beta,D_{2};b,\alpha,D_{1})]|^{2}\rangle.$
When these four different paths are distinguishable, the two-photon probability function equals
@@
$=$ $\frac{1}{4}[\langle|A(a,\alpha,D_{1};b,\beta,D_{2})|^{2}\rangle+\langle|A(a,\alpha,D_{2};b,\beta,D_{1})|^{2}\rangle$
$+\langle|A(a,\beta,D_{1};b,\alpha,D_{2})|^{2}\rangle+\langle|A(a,\beta,D_{2};b,\alpha,D_{1})|^{2}\rangle]$
Again, assume photons have equal probability being detected by D1 and D2. The ratio between Eqs. (IV.1.2) and (IV.1.2) at the same space-time coordinate equals the degree of second-order coherence,
$g^{(2)}(0)=p^{(2)}(0)=\frac{4|A^{\prime}{}|^{2}}{|A^{\prime}{}|^{2}}=4,$ (62)
@@
Figure 14: Scheme to observe two-photon superbunching with $2N$ different paths Zhou et al. (2017). P1, P2, …, PN are $N$ pinholes. RG1, RG2, …, RGN are $N$ rotating groundglasses.